# ViMo: A Generative Visual GUI World Model for App Agent

### v1

@@ 4 Experiments
In this section, we begin by summarising our proposed STR dataset discussed in Subsection 3.2.1. Next, we test the core capability of our ViMo, focusing on its GUI generation ability. Building on this, we demonstrate how the powerful GUI generation capability of ViMo can enhance the decision-making of App agents in real-world App operation tasks. Finally, we carry out the ablation study to validate the effectiveness of our model design. Specific setups and experiment details are elaborated in subsequent sections.
Table 1: Summarisation of our STR dataset.
App Episode Image Action
19 3550 23620 18450
@@ 4.1 Dataset Summarisation
Our STR dataset was constructed using data from two widely recognised and large-scale sources: Android Control Li et al. (2024) and Android in the Wild (AITW) Rawles et al. (2024b). From these sources, we respectively sampled 12 and 7 Apps, selecting those with rich data samples while filtering out noise. Android Control provides two types of user actions: (1) action commands: predefined actions (e.g., click, scroll) accompanied by specific parameters such as coordinates (x, y); (2) action instructions: actions described in natural languages, such as “click the plus icon”. We used action instructions as conditions for our world model as this approach was more concrete and better utilised the pre-trained model in understanding natural language. For AITW, action commands were converted into action instructions using OpenAI’s ChatGPT-4o. Our data is summarised in Table 1. To ensure both time-efficient and cost-efficient experiments, we followed prior App agent Rawles et al. (2024b) on partial split evaluation, instead of evaluating the model on the entire testing set. Specifically, we randomly sampled 57 episodes across 19 distinct Apps. More details about our dataset including collection details, split summarisation and full-split experiments are provided in the Appendix.
Figure 4: GUI generation comparison in graphic generation (Top) and text generation (Bottom, cropped for space efficiency).
@@ 4.2 World Model Ability
We evaluated the GUI generation capability of ViMo, from two critical perspectives: GUI quality evaluation and GUI-based trajectory synthesis. We compared against two types of world models: language-based and vision-based models. For language-based models, we adapted website world models Gu et al. (2024) to App setups, utilising Change-text to generate textual descriptions capturing differences between consecutive observations and HTML-text to predict App observations in an HTML format. For vision-based GUI world models, we included InstructPix2Pix (IP2P) Brooks et al. (2023) and UI-diffuser Wei et al. (2024), both originally designed for image editing and GUI generation. Further, we leveraged HTML-vision, where HTML-text were converted into images for visual representation.
Table 2: GUI quality evaluation. $s_{gc}$ indicates the GUI consistency, $s_{ia}$ instructional accuracy and $s_{ar}$ action readiness score. $s_{h}$ indicates the harmonic average between the 3 metrics. $\Delta s_{h}$ is the relative performance gains of our ViMo over other methods.
Method $s_{gc}$ $s_{ia}$ $s_{ar}$ $s_{h}$ $\Delta s_{h}$
HTML-vision 0.6969 85.77 62.79 0.7154 5.39%
IP2P 0.7414 63.57 70.15 0.6901 10.20%
UI-diffuser 0.5977 39.61 38.75 0.4426 71.82%
ViMo (Ours) 0.7423 75.39 78.68 0.7605 -
Table 3: Trajectory synthesis evaluation. “T+L” denotes the accuracy of the whole trajectory with length L.
World Model T+1 T+2 T+3 T+4
w/o world model 22.81 14.04 7.02 0
Change-text 52.63 26.32 10.53 5.26
HTML-text 38.60 14.04 12.28 7.02
HTML-vision 43.86 19.30 10.53 10.53
UI-diffuser 52.63 29.82 12.28 5.26
IP2P 56.14 21.05 10.53 7.02
ViMo (Ours) 57.89 36.84 14.03 12.28
@@
GUI Quality Evaluation. In this setting, we compared our approach against the three visual world models: IP2P, UI-diffuser and HTML-vision. For the evaluation metrics: The GUI consistency score ($s_{gc}$) assessed the visual similarity between the ground truth and the generated next GUI using DINO Caron et al. (2021) as the visual encoder; Instructional accuracy score ($s_{ia}$) used an LLM to determine whether the generated GUI adheres to the user action; Action readiness score ($s_{ar}$) evaluated whether the generated GUI retains valid elements essential for subsequent actions required to achieve the user goal. Prompts for $s_{ia}$ and $s_{ar}$ are given in the Appendix.
@@
As shown in Table 2, ViMo achieved the highest score on the harmonic average of the three metrics, surpassing other methods with an average relative performance improvement of 29.14%. Qualitative comparisons are presented in Fig. 4, under two scenarios: GUI graphic changes (Top) and text generation (Bottom). Experiments revealed that while the HTML-vision method exhibited greater flexibility in responding to user actions (see the bottom of Fig. 4 and $s_{ia}$ in Table 2), it failed to produce concrete details necessary for future actions, as indicated by lower $s_{ar}$. Conversely, IP2P generated plausible GUI graphics but lacked flexibility in text content generation, (as reflected by $s_{ia}$ in Table 2). This trade-off highlighted the superior balance of ViMo.
Additionally, we studied the generalisation of ViMo in Fig. 5 by providing user actions that were not typically encountered within the App’s standard context. For example, in the Clock App, a user action to “add a file” generated a Drive-style file selection window while retaining the Clock interface. Similarly, in the Kitchen Store App, ViMo can generate content corresponding to the action. These results emphasised ViMo’s generalisation ability facing novel combinations of App observations and user actions.
GUI-based Trajectory Synthesis. In this experiment, ViMo generated GUIs served as the observation for App agents to predict an action to achieve the user goal, and the agent’s output actions were evaluated. This setup assessed whether ViMo provided concrete and reliable details that aligned with real-world scenarios for action prediction. The process was iterated for L steps and the success accuracy of the entire L-step trajectory was calculated. More details about the agent setup and evaluation are in the Appendix.
As shown in Table 3, we compared ViMo against both visual- and language-based world models and demonstrated that our model significantly outperformed the other methods by providing more accurate and reliable information. This was reflected in higher trajectory prediction accuracy, underscoring the ability of our model to generate GUIs that aligned with the real-world environment.
@@ 4.3 World Model Enhanced App Agent
Agent Type App Agent Step Acc.
Language-Based ER Li et al. (2024) 34.50
AutoDroid Wen et al. (2024a) 35.46
T3A Rawles et al. (2024a) 43.13
T3A + VisMo (Ours) 49.20
Multi-Modality-Based APP-Agent Zhang et al. (2023) 42.81
Mobile-Agent-v2 Wang et al. (2024a) 43.45
M3A Rawles et al. (2024a) 46.01
M3A + VisMo (Ours) 50.16
@@
Table 5: Decision optimisation comparisons on APP agent performance. Apps are categorised into “Leisure”, “Work”, and “System”.
@@
M3A w/o world model w/o world model 46.62 51.28 43.65 46.01
Langugae Change-text 51.35 51.28 41.27 47.28
HTML-text 50.68 51.28 40.48 46.65
Vision HTML-vision 52.03 48.72 45.24 48.89
UI-diffuser 50.00 48.72 44.44 47.60
IP2P 52.03 48.72 44.44 48.56
ViMo (Ours) 53.38 53.85 45.24 50.16
In this section, we demonstrate that: (1) ViMo enhanced the performance of App agents in decision-making; (2) ViMo outperformed other world models in enabling App agents to make more accurate decisions.
Comparison with App Agents. In this experiment, we collected 6 LLM-based App agents, which included three language-based methods: ER, AutoDroid, and T3A, as well as three multi-modality-based methods: APP-Agent, Mobile-Agent-v2, and M3A. We applied our ViMo into M3A and T3A following the process in Subsection 3.3. Moreover, we followed the previous works Rawles et al. (2024b); Li et al. (2024) to use the step accuracy (the number of correct actions divided by the number of overall actions) to quantify the model performance. Table 4 demonstrates that the ViMo was beneficial to the App agent, achieving a relative performance gain of 9.01% for M3A and 14.07% for T3A. These findings highlighted the effectiveness of our proposed world model in providing App agents with enhanced decision-making capability.
Comparison with World Models. We further evaluated the ability of ViMo to enhance App agent decision-making by comparing it against other world models discussed in Subsection 4.2. For more comprehensive comparisons, we applied all the world models on two App agents: language-based T3A and multi-modality-based M3A. To provide more detailed results, we categorised the Apps into three groups: “Leisure”, “Work” and “System”. Table 5 illustrates that our proposed world model outperformed other world models, achieving overall scores of 49.03% and 50.16% respectively with the two agents. These findings underscored the effectiveness and practical value of our world model in real-world applications. Additional information about the categorisation and experiments with more App agents are provided in the Appendix.
Figure 6: Ablation study on static text generation: (a) Generating static text using an LLM; (b) Retaining the text in the image and rendering it as image pixels.
Figure 7: STR generation with action command (a) and action instruction (b).
4.4 Ablation Study
In this section, we ablated on two key components of ViMo: (1) preserving static text within the image to simplify the text generation task, and (2) using action instructions instead of action commands as the conditioning input for ViMo.
Firstly, for the challenge of predicting static text from specific GUI elements, such as keyboard, number pad or clock face, which typically do not involve text changes and exhibit complex spatial patterns, we retained static text within the image (Subsection 3.2.1). This approach eliminates the need for the LLM to generate such static text while generating in pixels instead. Fig. 6 illustrates the difficulties faced by the LLM when attempting to predict the static text associated with complex spatial patterns.
Secondly, we proposed conditioning STR prediction on action instruction rather than action commands (Subsection 4.1). Fig. 7 displays the STR generation of the same user intent but with different action types. It demonstrated that models learned with action commands failed to predict STR that aligns with the user’s intent, whereas action instructions offered a more concrete description, enabling the model to better capture the intent. Table 6 quantitatively confirms that both components significantly enhanced the performance of our ViMo on two App agents (T3A and M3A), underscoring their critical role in generating high-quality GUIs.
Table 6: Ablation study for the effects of preserving static text and using action instructions on the performance of App agents. “Static Text” indicates whether the static text was preserved in images. “Action Instructions” denotes whether action instructions ($\surd$”) or action commands (“-”) were used as the conditioning to ViMo.
Static Text Action Instruction App Agent
T3A M3A
- - 43.13 46.01
$\surd$ - 42.81 45.05
- $\surd$ 47.28 48.88
$\surd$ $\surd$ 49.20 50.16
5 Conclusion
In this work, we introduced ViMo, a novel generative visual GUI world model designed to predict App observations in a visual modality, providing a more realistic and concrete approach compared to contemporary language-based models. To address the unique challenges of GUI generation, ViMo is equipped with the Symbolic Text Representation (STR) to simplify text content generation to text location prediction by overlaying text content with placeholders and delegating content generation to LLM. This innovation ensured high visual fidelity and avoided artefacts like distorted or blurred text. Through extensive experiments, we demonstrated that ViMo generated both visually plausible and functionally effective GUIs. Notably, ViMo boosted step-wise action prediction accuracy by a relative performance gain of 14.07%, underscoring its potential to enhance App agents’ decision-making capabilities. While we required data collection for STR predictor training, we displayed a generalisation ability in generating GUIs conditioned on a novel combination of the current GUI and an action. These results highlighted the value of ViMo for advancing GUI world models and set a new benchmark for future research.
Appendix A GUI-text Predictor Details
This section elaborates on the design and functionality of the GUI-text predictor, summarising its key components and providing a detailed explanation of its underlying processes.
The GUI-text predictor operates by locating the text symbols. To be specific, we first detect black borders by identifying black pixels in the BGR colour space, generating a binary mask that indicates whether a pixel is black or not. A pixel is classified as black if its BGR values fall within the range [0,0,0] to [50,50,50]. Next, we identify rectangular regions within this mask by computing the ratio of the actual contour area to its corresponding bounding rectangle area. If this ratio exceeds 0.8, the region is considered a valid rectangle, allowing us to extract rectangles with black borders. For these detected regions, we further analyse their internal colour distribution to determine whether they contain the desired white colour. Specifically, we define white pixels as those with BGR values within the range [200,200,200] to [255,255,255]. If more than 50% of the pixels within a region fall within this range, the region is classified as a text symbol. Thus, the locations of text symbols are extracted, and we assign a unique identifier (ID) to each symbol through enumeration.
Building on this, we take as inputs the current GUI image $x_{k}$, an action $a$ to be applied to this image, the predicted STR ( $\text{STR}_{x_{k+1}^{a}}$), the location and unique ID token of the text symbols in the STR $\mathcal{T}$. Then we leverage an LLM to predict the text content for each text symbol.
The process begins with preprocessing the STR by overlaying the ID token for each text symbol to the corresponding position in the STR image, resulting in a modified representation denoted as $\text{STR}^{ID}_{k+1}$. Next, we prompt an LLM to identify which text symbols will remain unchanged after the action $a$ (see the prompt in Subsection G.1). These symbols are determined to not be affected by the action and have content identical to the previous GUI $x_{k}$. Based on the resulting ID list, we retrieve the corresponding pixels from the previous GUI $x_{k}$ based on their location and update the STR representation. The updated image is still referred to as $\text{STR}^{ID}_{k+1}$ for simplicity.
Finally, to overlay a symbol with its actual text content, we perform the following steps: (1) For a given text symbol’s location and corresponding text, the average background colour is computed by the average colour of the area on the edge of text symbol’s coordinates; (2) The text colour is set to either white or black to ensure optimal contrast with the background colour, for better visibility; (3) The font size is calculated as the maximum size that allows the text to fit entirely within the boundaries of the text symbol, ensuring optimal use of space and readability.
Table 7: Summarisation of our STR dataset.
Split App Episode Image Instrucion
Train 19 2853 19010 14852
Val 19 349 2290 1774
Test 19 348 2320 1824
All 19 3550 23620 18450
Appendix B Action Selection Details
Table 8: Decision optimisation comparisons on APP agent performance. Apps are categorised into “Leisure”, “Work”, and “System”.
@@ Appendix B Action Selection Details
Appendix C Data Collection
To ensure the quality and diversity of data samples for each App, while minimising noise, we collected App information from both Android Control Li et al. (2024) and Android in the Wild dataset (AITW) Rawles et al. (2024b) datasets. To be specific, out of 15,274 episodes in the Android Control, only 5,697 episodes include the “open_app” action. From these episodes, we extracted their “app_name”, identifying 758 unique applications. However, only 13 of these Apps had more than 50 samples. To enrich the dataset, we manually collected additional samples for these 13 Apps from the rest of the dataset. For AITW, we extracted App names by using the package name listed under the “current activity” field. After filtering out noisy, 11 valid Apps remained. By combining the overlapping applications from both datasets, we obtained a total of 19 unique Apps. We split our dataset into “Train”, “Validation” and “Test” splits, and we summarise our dataset under each split in Table 7.
Furthermore, we converted action commands into action instructions for AITW with specific prompts in Subsection G.6. We use Paddleocr Shi et al. (2016) for STR generation.
Appendix D Evaluation Details

### v2

@@ 4 Experiments
In this section, we begin by summarising our proposed STR dataset discussed in Subsection 3.2.1. Next, we tested the core capability of our ViMo, focusing on its GUI generation ability. Building on this, we demonstrated how the powerful GUI generation capability of ViMo can enhance the decision-making of App agents. Then, we studied our effectiveness in real-world App navigation tasks. Finally, we carried out the ablation study to validate the effectiveness of our model design. Specific setups and experiment details are elaborated in subsequent sections. Unless explicitly stated otherwise, GPT-4o [47] was employed as the default LLM in the following sections.
@@ 4.1 Dataset Summarisation
Our STR dataset was constructed using data from two widely recognised and large-scale sources: Android Control [21] and Android in the Wild (AITW) [4]. From these sources, we respectively sampled 12 and 7 Apps, selecting those with rich data samples while filtering out noise. Android Control provides two types of user actions: 1) action commands: predefined actions (e.g., click, scroll) accompanied by specific parameters such as coordinates (x, y); 2) action instructions: actions described in natural languages, such as "click the plus icon". We used action instructions as conditions for our world model as this approach was more concrete and better utilised the pre-trained model in understanding natural language. For AITW, action commands were converted into action instructions using GPT-4o [47]. In total, we collected 19 Apps with 3,550 episodes, 23,620 images and 18,450 actions. To ensure both time-efficient and cost-efficient experiments, we followed prior App agent [4] on partial split evaluation, instead of evaluating the model on the entire testing set. Specifically, we randomly sampled 57 episodes across 19 distinct Apps. Details including dataset collection, split summarisation and full-split experiments are provided in the Appendix.
Figure 3: GUI generation comparison in graphic generation (Top) and text generation (Bottom).
@@ 4.2 World Model Ability
We evaluated the GUI generation capability of ViMo by GUI quality evaluation. We included IP2P [11] and UI-diffuser [9], both originally designed for image editing and GUI generation. For IP2P, we fine-tuned their diffusion model on our dataset to generate everything of the GUI, including the text content and the graphic, denoted as IP2P*. We also leveraged an LLM to predict App observations in an HTML format, which were rendered into images, denoted as HTML-vision.
@@
We leveraged 3 evaluation metrics: The GUI consistency score ($s_{gc}$) assessed the visual similarity between the ground truth and the generated next GUI; Instructional accuracy score ($s_{ia}$) determined whether the generated GUI adheres to the user action; Action readiness score ($s_{ar}$) evaluated whether the generated GUI retains valid elements essential for subsequent actions required to achieve the user goal. We conducted both automatic evaluations and user preference studies to assess our methods. For the automatic evaluation, we used DINO [48] as the visual encoder to compute $s_{gc}$, and an LLM to evaluate $s_{ia}$ and $s_{ar}$. The prompts used for the LLM-based evaluations are provided in the Appendix. For the user study, we invited 70 voluntary participants to complete questionnaires based on 80 GUI samples, generated by all 4 compared methods. For each sample, participants were asked three questions—one each for evaluating $s_{gc}$, $s_{ia}$, and $s_{ar}$. The full instructions provided to the participants are provided in the Appendix.
As shown in Table 1, ViMo achieved the highest score on the harmonic average of the three automatic metrics, surpassing other methods with an average relative performance improvement of 29.14%. The results of the user study were consistent with the automatic evaluations, where ViMo demonstrated the best performance. Notably, HTML-vision and UI-diffuser performed significantly worse in human evaluation compared to their scores in the LLM-based assessment. This discrepancy likely arose because human evaluators perceived the outputs of these methods as visually unrealistic or functionally incoherent, leading to lower subjective scores in $s_{gc}$, $s_{ia}$ and $s_{h}$.
@@
Qualitative comparisons are presented in Fig. 3, under two scenarios: GUI graphic changes (Top) and text generation (Bottom, cropped for space efficiency). Experiments revealed that while the HTML-vision method exhibited greater flexibility in responding to user actions (as shown in the bottom examples), it failed to produce concrete details necessary for future actions (top). Conversely, IP2P* generated plausible GUI graphics but lacked flexibility in text content generation (also reflected by $s_{gc}$ and $s_{ia}$ in Table 1). This trade-off highlighted the superior balance of ViMo.
@@ 4.3 World Model Enhanced App Agent
Agent Type App Agent Leisure Work System Overall
Language-Based ER [21] 31.76 46.15 34.13 34.50
AutoDroid [20] 35.81 46.15 31.75 35.46
T3A [5] 41.22 51.28 42.86 43.13
T3A + ViMo (Ours) 50.00 58.97 45.24 49.20
Multi-Modality-Based APP-Agent [49] 43.24 51.28 39.68 42.81
Mobile-Agent-v2 [6] 43.92 53.85 39.68 43.45
M3A [5] 46.62 51.28 43.65 46.01
M3A + ViMo (Ours) 53.38 53.85 45.24 50.16
This section demonstrates that: 1) ViMo enhanced the performance of App agents in decision-making; 2) ViMo outperformed other world models in enabling App agents to make more accurate decisions.
Comparison with App Agents. In this experiment, we collected 6 LLM-based App agents, which included three language-based methods: ER, AutoDroid, and T3A, as well as three multi-modality-based methods: APP-Agent, Mobile-Agent-v2, and M3A. We applied our ViMo into M3A and T3A following the process in Subsection 3.3. Moreover, we followed the previous works [4, 21] to use the step accuracy (the number of correct actions divided by the number of overall actions) to quantify the model performance. To provide more detailed results, we categorised the Apps into three groups: "Leisure", "Work" and "System". Table 2 demonstrates that ViMo was beneficial to the App agent, achieving a relative performance gain of 9.01% for M3A and 14.07% for T3A. These findings highlighted the effectiveness of our proposed world model in providing App agents with enhanced decision-making capability. Additional information about the categorisation and experiments with more App agents are provided in the Appendix.
Comparison with World Models. We further evaluated the ability of ViMo to enhance App agent decision-making by comparing it against existing world models. In addition to vision-based world models discussed in Subsection 4.2, we also incorporated two language-based world models where we adapted website world models [8] to App setups, utilising Change-text to generate textual descriptions capturing differences between consecutive observations and HTML-text to predict App observations in an HTML format. Then, we applied all the world models on M3A App agents. Table 5 illustrates that our proposed world model outperformed other world models, achieving a step accuracy (Step Acc.) of 50.16%. These findings underscored the effectiveness of our world model in enhancing step-level decision-making. Experiments with more App agents are provided in the Appendix.
Table 3: Result on M3A Agent.
Modality World Model Step Acc.
w/o WM w/o WM 46.01
Language Change-text 47.28
HTML-text 46.65
Vision HTML-vision 48.89
UI-diffuser 47.60
IP2P* 48.56
ViMo (Ours) 50.16
Table 4: Zero-shot Evaluation.
App Agent LLM Step Acc.
SeeAct GPT-4-Turbo 33.9
M3A GPT-4-Turbo 42.1
ER Gemini 1.5 Pro 24.4
T3A Gemini-2.0-Flash 41.4
T3A+ViMo Gemini-2.0-Flash 46.8
M3A Gemini-2.0-Flash 44.2
M3A+ViMo Gemini-2.0-Flash 47.6
Table 5: Online Evaluation.
App Agent LLM Task Acc.
SeeAct GPT-4-Turbo 15.50
M3A GPT-4-Turbo 25.40
M3A Gemini-1.5-Pro 22.80
T3A GPT-4-Turbo 30.60
T3A Gemini-1.5-Pro 19.40
T3A Gemini-2.0-Flash 33.19
T3A + ViMo Gemini-2.0-Flash 40.95
4.4 Real-world Applications
Practical Deployment.
ViMo was designed to be lightweight and easily deployable. The minimum requirement for deployment is a GPU with 16 GB of memory. Moreover, ViMo was implemented as a plug-and-play API that required only a single function call, making integration straightforward. Inference time on V100 GPU is 8 seconds on a STR image generation and 30 seconds on GUI-text prediction. We collected and compared the inference time with existing methods in the Appendix.
Generalisation to New Apps.
Generalisation is a crucial capability for real-world applications. To assess the generalisation performance of our method on new Apps that were unseen during training, we conducted a zero-shot evaluation using data from the Android Control dataset [21], explicitly excluding Apps encountered during training. As shown in Table 4, our approach substantially outperformed the baseline and achieved 47.6%, underscoring its robustness and adaptability to novel App environments. Additional visualisations of unseen scenarios are provided in the Appendix.
Online Evaluation.
To further demonstrate the effectiveness of ViMo in realistic App navigation scenarios, we conducted an online evaluation using the AndroidWorld dataset [5], which comprises 116 distinct navigation tasks. Performance was measured using the task success rate (Task Acc.). As illustrated in Table 5, ViMo achieved a notable improvement of 7.76% over the baseline method, highlighting its effectiveness and reliability in real-world settings.
4.5 Ablation Study
Table 6: Ablations on preserving static text and using action instructions.
Static Text Action Instr. App Agent
T3A M3A
N/A N/A 43.13 46.01
42.81 45.05
47.28 48.88
49.20 50.16
Table 7: Ablation on the number of iterations.
Method Iterations Step Acc. (%)
T3A N/A 39.94
T3A+ViMo 1 46.06
2 46.65
3 45.05
In this section, we ablated on three key components of ViMo: 1) preserving static text within the image to simplify the text generation task; 2) using action instructions instead of action commands as the conditioning input for ViMo; and 3) varying the number of iterations, where each iteration corresponds to one roll-out step into the future during GUI prediction.
Firstly, for the challenge of predicting static text from specific GUI elements, such as keyboard, number pad or clock face, which typically did not involve text changes and exhibited complex spatial patterns, we retained static text within the image (Subsection 3.2.1). This approach eliminated the need for the LLM to generate such static text while generating in pixels instead. Secondly, we proposed conditioning STR prediction on action instruction rather than action commands (Subsection 4.1). Ablation results are presented in Table 6, where "Static Text" indicates whether static text was retained in the images, and "Action Instr." denotes whether natural language instructions ($\surd$") or abstract action commands ("-") were used as conditioning input to ViMo. The first row indicates the baseline where ViMo was not applied. The table shows that both components contributed significantly to performance improvements across the two App agents, highlighting their critical roles in enabling ViMo to generate high-quality GUIs. Visual comparison examples are provided in the Appendix for further illustration.
Our ViMo predicted future GUI observations, which could be recursively fed back as input to simulate further into the future. In this ablation study, we varied the iteration number to evaluate how extended roll-outs impact prediction accuracy. We took Gemini-2.0-Flash [50] as the LLM in this study. As shown in Table7, performing two iterations yielded the highest accuracy. However, this also led to increased computational cost. Therefore, we selected one step as a practical trade-off between performance and efficiency. We also observed a slight decline in performance at iteration 3 relative to iterations 1 and 2, indicating that extending the prediction horizon did not necessarily improve agent behaviour. This was likely due to that longer horizons introduced not only additional foresight but also a greater accumulation of prediction errors, whose detrimental effect could outweigh the potential benefits. Further analysis of error accumulation across iterations, along with comparisons with various world models and implementation details, is provided in the Appendix.
5 Conclusion
In this work, we introduced ViMo, a novel generative visual GUI world model designed to predict App observations in a visual modality, providing a more realistic and concrete approach compared to contemporary language-based models. To address the unique challenges of GUI generation, ViMo was equipped with the Symbolic Text Representation (STR) to simplify text content generation to text location prediction by overlaying text content with placeholders and delegating content generation to LLM. This innovation ensured high visual fidelity and avoided artefacts like distorted or blurred text. Through extensive experiments, we demonstrated that ViMo generated both visually plausible and functionally effective GUIs. Notably, ViMo boosted step-wise action prediction accuracy by a relative performance gain of 14.07%, underscoring its potential to enhance decision-making of App agents. Furthermore, real-world experiments demonstrated the strong generalisation ability of ViMo to unseen Apps, along with its robust performance in online navigation tasks under real-time environment interaction. These results highlighted the value of ViMo for advancing GUI world models and set a new benchmark for future research.
Appendix A Experimental Details
A.1 GUI-text Predictor
This subsection elaborates on the design and functionality of the GUI-text predictor, summarising its key components and providing a detailed explanation of its underlying processes.
Given a STR prediction, the GUI-text predictor starts by locating the text symbols. To be specific, we first detect black borders by identifying black pixels in the BGR colour space, generating a binary mask that indicates whether a pixel is black or not. A pixel is classified as black if its BGR values fall within the range [0,0,0] to [50,50,50]. Next, we identify rectangular regions within this mask by computing the ratio of the actual contour area to its corresponding bounding rectangle area. If this ratio exceeds 0.8, the region is considered a valid rectangle, allowing us to extract rectangles with black borders. For these detected regions, we further analyse their internal colour distribution to determine whether they contain the desired white colour. Specifically, we define white pixels as those with BGR values within the range [200,200,200] to [255,255,255]. If more than 50% of the pixels within a region fall within this range, the region is classified as a text symbol. Thus, the locations of text symbols are extracted, and we assign a unique identifier (ID) to each symbol through enumeration.
Building on this, we take as inputs the current GUI image $x_{k}$, an action $a$ to be applied to this image, the predicted STR ( $\text{STR}_{x_{k+1}^{a}}$), the location and unique ID token of the text symbols in the STR $\mathcal{T}$. Then we leverage an LLM to predict the text content for each text symbol. The process begins with preprocessing the STR by overlaying the ID token for each text symbol to the corresponding position in the STR image, resulting in a modified representation denoted as $\text{STR}^{ID}_{k+1}$. Next, we prompt an LLM to identify which text symbols will remain unchanged after the action $a$ (see the prompt in Subsection G.1). These symbols are determined to not be affected by the action and have content identical to the previous GUI $x_{k}$. Based on the resulting ID list, we retrieve the corresponding pixels from the previous GUI $x_{k}$ based on their location and update the STR representation. The updated image is still referred to as $\text{STR}^{ID}_{k+1}$ for simplicity.
Finally, to overlay a symbol with its actual text content, we perform the following steps: 1) For a given text symbol’s location and corresponding text, the average background colour is computed by the average colour of the area on the edge of text symbol’s coordinates; 2) The text colour is set to either white or black to ensure optimal contrast with the background colour, for better visibility; 3) The font size is calculated as the maximum size that allows the text to fit entirely within the boundaries of the text symbol, ensuring optimal use of space and readability.
A.2 Action Selection
Table 8: Summarisation of our STR dataset.
        Split         App         Episode         Image         Instrucion
        Train         19         2853         19010         14852
        Val         19         349         2290         1774
        Test         19         348         2320         1824
        All         19         3550         23620         18450
A.3 Data Collection
To ensure the quality and diversity of data samples for each App, while minimising noise, we collected App information from both Android Control [21] and Android in the Wild dataset (AITW) [4] datasets. To be specific, out of 15,274 episodes in the Android Control, only 5,697 episodes include the "open_app" action. From these episodes, we extracted their "app_name", identifying 758 unique applications. However, only 13 of these Apps had more than 50 samples. To enrich the dataset, we manually collected additional samples for these 13 Apps from the rest of the dataset. For AITW, we extracted App names by using the package name listed under the "current activity" field. After filtering out the noisy, 11 valid Apps remained. By combining the overlapping applications from both datasets, we obtained a total of 19 unique Apps. We split our dataset into "Train", "Validation" and "Test" splits, and we summarise our dataset under each split in Table 8.
Furthermore, we converted action commands into action instructions for AITW with specific prompts in Subsection G.6. We use Paddleocr [43] for STR generation.
@@ A.3 Data Collection
Table 9: Decision optimisation comparisons on APP agent performance. Apps are categorised into "Leisure", "Work", and "System".
@@
M3A w/o world model w/o world model 46.62 51.28 43.65 46.01
Langugae Change-text 51.35 51.28 41.27 47.28
HTML-text 50.68 51.28 40.48 46.65
Vision HTML-vision 52.03 48.72 45.24 48.89
UI-diffuser 50.00 48.72 44.44 47.60
IP2P 52.03 48.72 44.44 48.56
ViMo (Ours) 53.38 53.85 45.24 50.16
A.4 Evaluation
World Model Ability.